### v1

@@ Abstract
Open-Vocabulary Mobile Manipulation (OVMM) is the problem of picking any object in any unseen environment, and placing it in a commanded location. This is a foundational challenge for robots to be useful assistants in human environments, because it involves tackling sub-problems from across robotics: perception, language understanding, navigation, and manipulation are all essential to OVMM. In addition, integration of the solutions to these sub-problems poses its own substantial challenges. To drive research in this area, we introduce the HomeRobot OVMM benchmark, where an agent navigates household environments to grasp novel objects and place them on target receptacles. HomeRobot has two components: a simulation component, which uses a large and diverse curated object set in new, high-quality multi-room home environments; and a real-world component, providing a software stack for the low-cost Hello Robot Stretch to encourage replication of real-world experiments across labs. We implement both reinforcement learning and heuristic (model-based) baselines and show evidence of sim-to-real transfer. Our baselines achieve a 20% success rate in the real world; our experiments identify ways future research work improve performance. See videos on our website: https://ovmm.github.io/.
@@ 1 Introduction
OVMM: We propose the first reproducible mobile-manipulation benchmark for the real world, with an associated simulation component. In simulation, we use a dataset of $200$ human-authored interactive 3D scenes [19] instantiated in the AI Habitat simulator [20, 21] to create a large number of challenging, multi-room OVMM problems with a wide variety of objects curated from a variety of sources. Some of these objects’ categories have been seen during training; others have not. In the real world, we create an equivalent benchmark, also with a mix of seen and unseen object categories, in a controlled apartment environment. We use the Hello Robot Stretch [22]: an affordable and compliant platform for household and social robotics that is already in use at over 40 universities and industry research labs. Fig. 1 shows instantiations of our OVMM task in both the real-world benchmark and in simulation. We have a controlled real-world test environment, and plan to run the real world benchmark yearly to assess progress on this challenging problem. Real-world benchmarking will be run as a part of the NeurIPS 2023 HomeRobot OVMM competition [23].
@@
HomeRobot: We also propose HomeRobot, [^1] https://github.com/facebookresearch/home-robot a software framework to facilitate extensive benchmarking in both simulated and physical environments. It comprises identical APIs that are implemented across both settings, enabling researchers to conduct experiments that can be replicated in both simulated and real-world environments. Table 3 compares OVMM+HomeRobot to the literature. The HomeRobot library also supports a number of sub-tasks, including manipulation learning [24], navigation [25], and object-goal navigation [2].
In this paper, we use HomeRobot to compare two families of approaches: a heuristic solution, using a motion planner shown to work for real-world object search [2], and a reinforcement learning (RL) solution, which learns how to navigate to objects given depth and predicted object segmentation. We use the open-vocabulary object detector DETIC [26] to provide object segmentation for both the heuristic and RL policies. We observe that while the RL methods moved to the object more efficiently if an object was visible, the heuristic planner was better at long-horizon exploration. We also see a substantial drop in performance in switching from from ground-truth segmentation to DETIC segmentation. This highlights the importance of the OVMM challenge, as only through viewing the problem holistically - integrating perception, planning, and action - can we build general-purpose home assistants.
@@
Table 1: Comparisons of our proposed benchmark with prior work. We provide a large number of environments with a continuous action space, and uniquely provide a real-world robotics stack with demonstrated sim-to-real capabilities, allowing others to reproduce and deploy their own solutions. Additional nuances in footnote [^3] ALFRED uses object masks for interaction. ObjectNav uses scans, not full object meshes. ProcThor scenes are procedurally generated, this has the benefit that the potential number of environments is unbounded.. Partial availability Not available Capability available
@@ 2 Related Work
Reproducibility of robotics research. Standardized robotics benchmarks have been pursued for a long time, often by open-sourcing robot designs or introducing low-cost robots [40, 41, 42, 43, 44, 45, 46, 47, 48]. However, the environments in which these robots are used vary dramatically, leading to evaluation of components (e.g., object navigation, SLAM) in isolation, instead of as components of a larger system that may not benefit from those changes. The HomeRobot stack enables end-to-end benchmarking of individual components by providing a full robotics stack, with multiple implementations of different sub-modules. The simplicity helps move beyond standardized sets of objects (e.g., [49, 50, 51]) to a common set of robots, objects, and environments.
@@ 3 Open-Vocabulary Mobile Manipulation
Crucially, we need and develop both (1) a simulation version of this OVMM problem, for reproducibility, training, and fast iteration, and (2) a real-robot stack with a corresponding real-world benchmark. We compare the two in Fig. 2. Our simulated environments allow for varied, long-horizon task experimentation; our real-world HomeRobot stack allows for experimenting with real data, and we design a set of real-world tests to evaluate the performance of our learned and heuristic baselines.
The Robot. We use the Hello Robot Stretch [22] with DexWrist as the mobile manipulation platform, because it (1) is relatively affordable at \$$25,000$ USD, (2) offers 6 DoF manipulation, and (3) is human safe and human-sized, making it safe to test in labs [24, 11] and homes [2], and can reach most places a human would expect a robot to go. For a breakdown on hardware choices, see Sec. G.1.
@@
Receptacles. We include common household receptables (e.g. tables, chairs, sofas) in our dataset; unlike with manipulable objects, all possible receptacle categories are seen during training.
Scenes. We have both a simulated scene dataset, and a fixed set of real-world scenes with specific furniture arrangements and objects. In both simulated and real scenes, we use a mixture of objects from previously-seen categories, and objects from unseen categories as the goal object for our Open-Vocabulary Mobile Manipulation task. We hold out validation and test scenes, which do not appear in the training data; while some receptacles may re-appear, they will be at previously-unseen locations, and target object instances will be unseen.
Scoring. We compute success for each stage: finding object on start_receptacle, successfully picking up object, finding goal_receptacle, and placing object on the goal. Overall success is true if all four stages were accomplished. We also compute a single partial success metric as a tie-breaker, in which agents receive 1 point for each successive stage accomplished per episode, normalized by the number of stages. More details in Appendix B.
@@ 3.1 Simulation Dataset
The Habitat Synthetic Scenes Dataset (HSSD) [19] consists of 200+ human authored 3D home scenes containing over 18k individual models of real-world objects. Like most real houses, these scenes are cluttered with furniture and other objects placed into realistic architectural layouts, making navigation and manipulation similarly difficult to the real world. We used a subset of HSSD [19] consisting of $60$ scenes (e.g. Fig. 4) for which additional metadata and simulation structures were authored to support rearrangement  [^4] All 200+ scenes with rearrangement support will be released at the time of final submission.. For our experiments these are divided into train, validation, and test splits of 38, 12, and 10 scenes each, following the splits in the original HSSD paper [19].
Figure 4: Example (object free) top-down view from HSSD [19]. See App. Figs. 8 & 10 for navigation and viewpoints.
Objects and Receptacles. We aggregate objects from AI2-Thor [57], Amazon-Berkeley Objects [58], Google Scanned Objects [59] and the HSSD [19] dataset to create a large and diverse dataset of real-world robot problems. In total, we annotated 2,535 objects from 129 total categories.We identified 21 different categories of receptacle which appear in the HSSD dataset [19].
@@
Table 2: # of objects in each split: (S)een, (U)nseen, (I)nstance, and (C)ategory in simulation.
Episode Generation. We generate episodes consisting of varying object arrangements and particular values for object, start_receptacle, and goal_receptacle, which allow our agent to successfully move about and interact with the world. In the case of Open-Vocabulary Mobile Manipulation, this task is particularly challenging because we have to place objects in locations which are navigable, meaning that the robot can get to them, reachable, meaning its arm can make it to these locations, and from which we can navigate to a navigable, reachable goal receptacle. For full episode generation details see App. C.2.
@@ 3.2 Real World Benchmark
During real-world testing, we selected a pool of object instances that did not appear during simulation training, but split between classes that did and did not appear in training. We used eight different categories, of which five were seen during training (Cup, Bowl, Stuffed Toy, Medicine Bottle, and Toy Animal), and three were not (Rubik’s cube, Toy Drill, and Lemon). We performed $20$ experiments on the Stretch robot for each of our two different baselines and with seven different receptacle classes: Cabinet, Chair, Couch, Counter, Sink, Stool, Table.
@@ 4 The [Uncaptioned image]HomeRobot Library
Figure 5: Refer to captionHomeRobot is a simple, easy-to-set-up library which works in multiple environments and requires only relatively affordable hardware. Computationally intensive operations are performed on a desktop PC with a GPU, and a dedicated consumer-grade router provides a network interface to a robot running low-level control and SLAM.
To facilitate research on these challenging problems, we open-source the HomeRobot library, which implements navigation and manipulation capabilities supporting Hello Robot’s Stretch [22]. In our setup, it is assumed that users have access to a mobile manipulator and a NVIDIA GPU powered workstation. The mobile manipulator runs the low-level controller and the localization module, while the desktop runs the high-level perception and planning stack(Fig. 5). The robot and desktop are connected using an off-the-shelf router [^5] Our experiments used a NetGear Nighthawk router.. HomeRobot is designed as a user-friendly software stack, enabling quick setup of the robot for immediate testing. The key features of our stack include:
@@ 4.1 Baseline Agent Implementation
Crucially, we provide baselines and tools that enable researchers to effectively explore the Open-Vocabulary Mobile Manipulation task. We include two types of baselines in HomeRobot: a heuristic baseline, in which we use a well known motion planning technique [2] and simple rules to execute grasping and manipulation actions; and a reinforcement learning baseline, where we learn exploration and manipulation skills using an off-the-self policy learning algorithm, DDPPO [60]. In addition, we have implemented example projects from several recently released papers, testing different capabilities such as object-goal navigation [1, 2], skill learning [24], continual learning [61], and image instance navigation [25]. Due to the challenging, long-horizon nature of the task, we implement a high-level policy called OVMMAgent which calls a sequence of skills to accomplish a task. We breakdown our agents into four skills:
@@
Grasp: Pick up the object. We provide a high-level action for this, since we do not simulate the gripper interaction in Habitat. However, our library is compatible with a range of learned grasping skills and supports learning policies for grasping.
@@
Heuristic. We implement a version using only off-the-shelf learned models and heuristics, noting that previous work in mobile manipulation has used these models to great effect (e.g. [62]). Here, DETIC [63] provides masks for an open-vocabulary set of objects as appropriate for each skill. The start_receptacle, object,goal_receptacle for each episode is given. Fig. 14 shows an example of the heuristic navigation and place policy being executed in the real world (App. D).
@@ 5 Results
Simulation Results Skill Partial Success Rates Overall Success Rate Partial Success Metric Perception Navigation Gaze Place FindObj Pick FindRec Ground Truth Heuristic None Heuristic 046.2 039.5 018.6 005.9 027.5 Heuristic RL RL 047.2 041.7 027.1 011.0 031.7 RL None Heuristic 055.1 041.9 026.4 005.2 032.1 RL RL RL 055.7 050.2 035.2 011.6 038.2 DETIC [26] Heuristic None Heuristic 023.3 011.5 003.0 000.3 009.5 Heuristic RL RL 024.8 009.5 005.0 000.2 009.9 RL None Heuristic 019.9 010.2 004.4 000.6 008.8 RL RL RL 019.8 011.8 006.3 000.8 009.7
Table 3: We observe that one of the main causes of failures for our baseline systems was perception failures; ground-truth performance is notably higher. We also see that both RL and heuristic skills struggled navigating tightly constrained multi-room environments and successfully placing objects.
@@
We report results in Table 3 broken down by skill. The results show that RL policies outperformed heuristic methods for both navigation and placement tasks. However, all policies experienced a decline in performance when the perception is changed from ground truth to DETIC. Notably, heuristic policies exhibited less degradation in performance compared to RL policies under DETIC perception. With DETIC perception, the heuristic FindObj policy outperforms RL. We attribute this to the heuristic policy’s ability to incorporate noisy predictions by constructing a 2D semantic map, which proves advantageous in handling small objects that are prone to misclassification. Furthermore, we observed that the learned gaze policy generally led to improved pick performance, except when used in combination with the Heuristic nav with DETIC perception. For additional information, example simulation trajectories can be found in Appendix Figure 16, and results comparing seen versus unseen categories are discussed in Section F.2.
Real World FindObj Pick FindRec Overall Success
Heuristic Only 0.70 0.35 0.30 0.15
RL Only 0.70 0.45 0.30 0.20
Table 4: Results for heuristic and RL baseline in the real world on the OVMM task. In both cases, the grasping action is executed as described in Sec. 4; but initial conditions of the robot such as its position relative to the object or to other obstacles may cause various failures.
Real World. Finally, we conducted a series of experiments in a real-world held-out apartment setting. We performed a total of $20$ episodes, utilizing a combination of seen and unseen object classes as our target objects. The results of these experiments are presented in  Table 4. RL performed slightly better than the Heuristic baseline, successfully completing 1 extra episode and achieving a success rate of 20%. This difference primarily stemmed from the pick and place sub-tasks. In the pick task, the RL Gaze skill plays a crucial role in achieving better alignment between the agent and the target object, which led to more successful grasping. Similarly, the RL place skill demonstrated more precision, ensuring that the object stayed closer to the surface of the receptacle.
@@
6 Limitations
Our benchmark has a few key limitations: (1) Due to simulation limitations, we don’t physically simulate grasping in the first version, which is why we provide a separate policy for this in the real world. Grasping is a well-studied problem [64, 65, 66], but simulations that train useful real-world grasp systems require special consideration. (2) We consider full natural language queries out-of-scope. Finally, (3) we do not implement many motion planners in HomeRobot (see Sec. D.2), or task-and-motion-planning with replanning, as would be ideal [67].

### v2

@@ Abstract
Open-Vocabulary Mobile Manipulation (OVMM) is the problem of picking any object in any unseen environment, and placing it in a commanded location. This is a foundational challenge for robots to be useful assistants in human environments, because it involves tackling sub-problems from across robotics: perception, language understanding, navigation, and manipulation are all essential to OVMM. In addition, integration of the solutions to these sub-problems poses its own substantial challenges. To drive research in this area, we introduce the HomeRobot OVMM benchmark, where an agent navigates household environments to grasp novel objects and place them on target receptacles. HomeRobot has two components: a simulation component, which uses a large and diverse curated object set in new, high-quality multi-room home environments; and a real-world component, providing a software stack for the low-cost Hello Robot Stretch to encourage replication of real-world experiments across labs. We implement both reinforcement learning and heuristic (model-based) baselines and show evidence of sim-to-real transfer of the nav and place skills. Our baselines achieve a 20% success rate in the real world; our experiments identify ways future work can improve performance. See videos on our website: https://ovmm.github.io/.
@@ 1 Introduction
OVMM: We propose the first reproducible mobile-manipulation benchmark for the real world, with an associated simulation component. In simulation, we use a dataset of $200$ human-authored interactive 3D scenes [19] instantiated in the AI Habitat simulator [20, 21] to create a large number of challenging, multi-room OVMM problems with a wide variety of objects curated from a variety of sources. Some of these objects’ categories have been seen during training; others have not. In the real world, we create an equivalent benchmark, also with a mix of seen and unseen object categories, in a controlled apartment environment. We use the Hello Robot Stretch [22]: an affordable and compliant platform for household and social robotics that is already in use at over 40 universities and industry research labs. Fig. 1 shows instantiations of our OVMM task in both the real-world benchmark and in simulation. We have a controlled real-world test environment, and plan to run the real-world benchmark yearly to assess progress on this challenging problem. Real-world benchmarking will be run as a part of the NeurIPS 2023 HomeRobot OVMM competition [23].
@@
HomeRobot: We also propose HomeRobot, [^1] https://github.com/facebookresearch/home-robot a software framework to facilitate extensive benchmarking in both simulated and physical environments. It comprises identical APIs that are implemented across both settings, enabling researchers to conduct experiments that can be replicated in both simulated and real-world environments. Table 3 compares HomeRobot OVMM to the literature. Notably, HomeRobot provides a robotics stack for the Hello Robot Stretch which supports a range of capabilities in both simulation and the real world, and is not restricted to just the OVMM task. Our library also supports a number of sub-tasks, including manipulation learning [24], continuous learning [25], navigation [26], and object-goal navigation [2].
In this paper, we use HomeRobot to compare two families of approaches: a heuristic solution, using a motion planner shown to work for real-world object search [2], and a reinforcement learning (RL) solution, which learns how to navigate to objects given depth and predicted object segmentation. We use the open-vocabulary object detector DETIC [27] to provide object segmentation for both the heuristic and RL policies. We observe that while the RL methods moved to the object more efficiently if an object was visible, the heuristic planner was better at long-horizon exploration. We also see a substantial drop in performance when switching from ground-truth segmentation to DETIC segmentation. This highlights the importance of the HomeRobot OVMM challenge, as only through viewing the problem holistically - integrating perception, planning, and action - can we build general-purpose home assistants.
@@
Table 1: Comparisons of our proposed benchmark with prior work. We provide a large number of environments and unique objects, focusing on manipulable objects, with a continuous action space. Uniquely, we also provide a multi-purpose, real-world robotics stack, with demonstrated sim-to-real capabilities, allowing others to reproduce and deploy their own solutions. Additional nuances in footnote [^3] ALFRED uses object masks for interaction. ObjectNav uses scans, not full object meshes. ProcThor scenes are procedurally generated, this has the benefit that the potential number of environments is unbounded.. Partial availability Not available Capability available
@@ 2 Related Work
Reproducibility of robotics research. Standardized robotics benchmarks have been pursued for a long time, often by open-sourcing robot designs or introducing low-cost robots [41, 42, 43, 44, 45, 46, 47, 48, 49]. However, the environments in which these robots are used vary dramatically, leading to evaluation of components (e.g., object navigation, SLAM) in isolation, instead of as components of a larger system that may not benefit from those changes. The HomeRobot stack enables end-to-end benchmarking of individual components by providing a full robotics stack, with multiple implementations of different sub-modules. The simplicity helps move beyond standardized sets of objects (e.g., [50, 51, 52]) to a common set of robots, objects, and environments. Ours is the only benchmark to provide a broadly capable robotics stack for implementing and sharing robotics code; this is similar to projects like PyRobot [53], which doesn’t also provide a strong simulation benchmark.
@@ 3 Open-Vocabulary Mobile Manipulation
Crucially, we need and develop both (1) a simulation version of the Open-Vocabulary Mobile Manipulation problem, for reproducibility, training, and fast iteration, and (2) a real-robot stack with a corresponding real-world benchmark. We compare the two in Fig. 2. Our simulated environments allow for varied, long-horizon task experimentation; our real-world HomeRobot stack allows for experimenting with real data, and we design a set of real-world tests to evaluate the performance of our learned and heuristic baselines.
The Robot. We use the Hello Robot Stretch [22] with DexWrist as the mobile manipulation platform, because it (1) is relatively affordable at \$$25,000$ USD, (2) offers 6 DoF manipulation, and (3) is human safe and human-sized, making it safe to test in labs [24, 11] and homes [2], and can reach most places a human would expect a robot to go. For a breakdown of hardware choices, see Sec. H.1.
@@
Receptacles. We include common household receptacles (e.g. tables, chairs, sofas) in our dataset; unlike with manipulable objects, all possible receptacle categories are seen during training.
Scenes. We have both a simulated scene dataset and a fixed set of real-world scenes with specific furniture arrangements and objects. In both simulated and real scenes, we use a mixture of objects from previously-seen categories, and objects from unseen categories as the goal object for our Open-Vocabulary Mobile Manipulation task. We hold out validation and test scenes, which do not appear in the training data; while some receptacles may re-appear, they will be at previously unseen locations, and target object instances will be unseen.
Scoring. We compute success for each stage: finding object on start_receptacle, successfully picking up object, finding goal_receptacle, and placing object on the goal. Overall success is true if all four stages were accomplished. We compute partial success as a tie-breaker, in which agents receive 1 point for each successive stage accomplished, normalized by the number of stages. More details in Appendix C.
@@ 3.1 Simulation Dataset
The Habitat Synthetic Scenes Dataset (HSSD) [19] consists of 200+ human-authored 3D home scenes containing over 18k 3D models of real-world objects. Like most real houses, these scenes are cluttered with furniture and other objects placed into realistic architectural layouts, making navigation and manipulation similarly difficult to the real world. We used a subset of HSSD [19] consisting of $60$ scenes for which additional metadata and simulation structures were authored to support rearrangement  [^4] All 200+ scenes with rearrangement support will be released soon.. For our experiments, these are divided into train, validation, and test splits of 38, 12, and 10 scenes each, following the splits in the original HSSD paper [19].
Objects and Receptacles. We aggregate objects from AI2-Thor [59], Amazon-Berkeley Objects [60], Google Scanned Objects [61] and the HSSD [19] dataset to create a large and diverse dataset of real-world robot problems. In total, we annotated 2,535 objects from 129 total categories.We identified 21 different categories of receptacles which appear in the HSSD dataset [19].
@@
Table 2: # of objects in the sim for each split of (S)een and (U)nseen (I)nstance and (C)ategory.
Episode Generation. We generate episodes consisting of varying object arrangements and particular values for object, start_receptacle, and goal_receptacle, which allow our agent to successfully move about and interact with the world. In the case of Open-Vocabulary Mobile Manipulation, this task is particularly challenging because we have to place objects in locations that are navigable, meaning that the robot can get to them, reachable, meaning its arm can make it to these locations, and from which we can navigate to a navigable, reachable goal receptacle. For full episode generation details see App. D.2.
@@ 3.2 Real World Benchmark
During real-world testing, we selected object instances that did not appear in simulation training, split between classes that did and did not appear. We used eight different categories: five seen (Cup, Bowl, Stuffed Toy, Medicine Bottle, and Toy Animal), and three unseen (Rubik’s cube, Toy Drill, and Lemon). We performed $20$ experiments for each of our two different baselines and with seven different receptacle classes: Cabinet, Chair, Couch, Counter, Sink, Stool, Table.
@@ 4 The [Uncaptioned image]HomeRobot Library
Figure 4: Refer to captionHomeRobot is a simple, easy-to-set-up library that works in multiple environments and requires only relatively affordable hardware. Computationally intensive operations are performed on a desktop PC with a GPU, and a dedicated consumer-grade router provides a network interface to a robot running low-level control and SLAM.
To facilitate research on these challenging problems, we open-source the HomeRobot library, which implements navigation and manipulation capabilities supporting Hello Robot’s Stretch [22]. In our setup, it is assumed that users have access to a mobile manipulator and an NVIDIA GPU-powered workstation. The mobile manipulator runs the low-level controller and the localization module, while the desktop runs the high-level perception and planning stack(Fig. 4). The robot and desktop are connected using an off-the-shelf router [^5] Our experiments used a NetGear Nighthawk router.. The key features of our stack include:
@@ 4.1 Baseline Agent Implementation
Crucially, we provide baselines and tools that enable researchers to effectively explore the Open-Vocabulary Mobile Manipulation task. We include two types of baselines in HomeRobot: a heuristic baseline, using motion planning [2] and simple rules for manipulation; and a reinforcement learning baseline. In addition, we have implemented example projects from several recently released papers, testing different capabilities such as object-goal navigation [1, 2], skill learning [24], continual learning [25], and image instance navigation [26]. We implement a high-level policy called OVMMAgent which calls a sequence of skills one after the other. These skills are:
@@
Pick: Pick up the object. We provide a high-level action for this, since we do not simulate the gripper interaction in Habitat. However, our library is compatible with a range of learned grasping skills and supports learning policies for grasping.
@@
Specifically, OVMMAgent is a state-machine that calls FindObj, Gaze, Pick, FindRec, and Place in that order, where Pick is a grasping policy provided by the robot library in the real world. The other skills are created using the approaches given below:
Heuristic. We implement a version using only off-the-shelf learned models and heuristics, noting that previous work in mobile manipulation has used these models to great effect (e.g. [62]). Here, DETIC [63] provides masks for an open-vocabulary set of objects as appropriate for each skill. The start_receptacle, object,goal_receptacle for each episode is given. Fig. 16 shows an example of the heuristic navigation and place policy being executed in the real world (App. E).
@@ 5 Results
Simulation Results Skill Partial Success Rates Overall Success Rate Partial Success Metric Perception Navigation Gaze Place FindObj Pick FindRec Ground Truth Heuristic None Heuristic 054.1 048.5 031.5 005.1 034.8 Heuristic RL RL 056.5 051.5 042.3 013.2 040.9 RL None Heuristic 065.4 054.8 043.7 007.3 042.8 RL RL RL 066.6 061.1 050.9 014.8 048.3 DETIC [27] Heuristic None Heuristic 028.7 015.2 005.3 000.4 012.4 Heuristic RL RL 029.4 013.2 005.8 000.5 012.2 RL None Heuristic 021.9 011.5 006.0 000.6 010.0 RL RL RL 021.7 010.2 006.2 000.4 009.6
Table 3: Partial and overall success rate (SR) (in %) for different combinations of skills and perception systems. The partial SR for each skill is dependent on the previous skill’s SR. The partial SR for the place skill is the same as the overall SR. The partial success metric is calculated by averaging the 4 partial SRs. One of the main causes of failures for our baseline systems was perception; ground-truth perception is notably better. Both RL and heuristic skills struggled to navigate tightly constrained multi-room environments and successfully place objects.
@@
We report results on HomeRobot OVMMin Table 3. RL policies outperformed heuristic methods for both navigation and placement tasks. However, all policies declined in performance when using DETIC instead of ground truth segmentation. Heuristic policies exhibited less degradation in performance compared to RL policies: when using DETIC, the heuristic FindObj policy even outperforms RL. We attribute this to the heuristic policy’s ability to incorporate noisy predictions by constructing a 2D semantic map, which helps handle small objects that are prone to misclassification. Furthermore, using the learned gaze policy led to improved pick performance, except when used in combination with the Heuristic nav with DETIC perception. Example simulation trajectories can be found in Appendix Figure 18, and comparisons of seen versus unseen categories in Section G.2.
Real World. Finally, we conducted a series of experiments in a real-world held-out apartment setting. We performed a total of $20$ episodes, utilizing a combination of seen and unseen object classes as our target objects. The results of these experiments are presented in  Table 4. RL performed slightly better than the Heuristic baseline, successfully completing 1 extra episode and achieving a success rate of 20%. This difference primarily stemmed from the pick and place sub-tasks. In the pick task, the RL Gaze skill plays a crucial role in achieving better alignment between the agent and the target object, which leads to more successful grasping. Similarly, the RL place skill demonstrated more precision, ensuring that the object stayed closer to the surface of the receptacle.
Real World FindObj Pick FindRec Overall Success Heuristic Only 70 35 30 15 RL Only 70 45 30 20
Table 4: Success Rate (in %) for heuristic and RL baselines in the real world OVMM task. In both cases, the grasping action is executed as described in Sec. 4; but initial conditions of the robot such as its position relative to the object or to other obstacles may cause various failures.